OpenStudy (anonymous):

PLEASE HELP!! how do i find a1 for the given geometric series?? sn=48455, r=3.2, n=3????

OpenStudy (anonymous):

what is the formula for finding a1 in a geometric series???

OpenStudy (anonymous):

can you help me hero? i have been trying to get help forever!!

hero (hero):

Hang on a sec :D

OpenStudy (anonymous):

thank you!! :D

hero (hero):

I think the formula is \(a_n = a_{1}r^{(n-1)}\) And to get \(a_1\) you simply divide both sides by \(r^{n-1}\) : \(a_1 = \large\frac{a_n}{a^{n-1}}\)

hero (hero):

Whoops

hero (hero):

\(a_1 = \large\frac{a_n}{r^{n-1}}\)

OpenStudy (anonymous):

but how do i solve for an if i dont have a1?

hero (hero):

So all you have to do is plug the numbers in

hero (hero):

\(a_1 = \large\frac{48455}{3.2^{(3-1)}}\)

OpenStudy (anonymous):

thank you!! i get it now:)

hero (hero):

You might want to confirm that with someone else. Maybe @amistre64

OpenStudy (anonymous):

ok will do!

hero (hero):

I found another formula so I don't know which one it is

OpenStudy (amistre64):

\[S_n=a1*\frac{1-r^n}{1-r}\]

hero (hero):

Yeah, that's the other one I found

OpenStudy (amistre64):

i had it in the cupboard behind the peanut butter cookies

OpenStudy (anonymous):

sn=48455, r=3.2, n=3???? Sn probably means the Sum of the first n terms, not the nth term, which would be An.

hero (hero):

I should never have found the wrong formula.

OpenStudy (anonymous):

ok...so solve for An then?

hero (hero):

The correct formula doesn't seem to have A_n in it

hero (hero):

Look at what @amistre64 posted.

OpenStudy (anonymous):

i already know the formula for sn...i need to solve for a1..which is what is confusing me so much

OpenStudy (anonymous):

Alright, so let's first just define some basic terms, okay? r= the common ratio \(\large A_n = \text{the nth term}\) \(\large S_n = \text{the sum of the first n terms}\) Now, the cool thing is that there is a formula for the sum of the first n terms of a geometric sequence. We'll plug into this formula everything that we know, and that will let us solve for \(a_1\).

OpenStudy (anonymous):

The formula we have is: \(\Large S_n = a_1*\frac{1-r^n}{1-r}\) The information we have is \(S_n=48455, r=3.2, n=3\) So let's plug that information into the formula we have.

OpenStudy (amistre64):

some people swap out a negtaive and write (r^n-1)/(r-1)

OpenStudy (anonymous):

\(\Large S_3 = 48455 = a_1*\frac{1-3.2^3}{1-3.2}\) Do you see how I get that? Ask me questions.

OpenStudy (anonymous):

okay..so i did that and got -1385.69...correct?..or no good

OpenStudy (anonymous):

ok..nvrm ignore my answer..i didnt put the 3 in for sn....continue

OpenStudy (anonymous):

You can substitute the 3 in wherever you have n, but that doesn't mean you substitute \(S_n \rightarrow 3\) Instead, you can substitute and get \(S_n \rightarrow S_3\)

OpenStudy (amistre64):

\[S=a*\frac tb\]\[\frac bt(S=a*\frac tb)\to\ S\frac bt = a\]

OpenStudy (anonymous):

So I'm not sure what you were trying there, but let's step back and make sure we understand the process. Step 1: Start with the correct general formula. In this case, it's \(\Large S_n = a_1 *\frac{1-r^n}{1-r}\) Step 2: Look at the information we're given and substitute it into the formula, leaving the unknown info as a variable. Step 3: Do some basic algebra, solve for the unknown. In this case that unknown is \(a_1\).

OpenStudy (anonymous):

So let's take this step by step. Do you understand step 1? Does that formula make sense to you, and do you know what all of the variables mean?

OpenStudy (anonymous):

yes looks good so far

OpenStudy (anonymous):

Great, step 2. Can you substitute in the things you know?

hero (hero):

I don't mean to interrupt. I've clearly posted the wrong formulas and things but @mathdummy4, now that you have the right formulas, all you have to do is isolate a_1 since it is what you need to find: \(a_1 = \large\frac{s^n}{\frac{1-r^n}{1-r}}\)

OpenStudy (anonymous):

We're getting there, hero.

hero (hero):

That's a shortcut @SmoothMath

OpenStudy (anonymous):

When you're working with a student, you have to go at their pace and make sure they understand each step. I know you and I are good with the shortcut, but I'd rather take a steady step by step approach for her sake.

OpenStudy (anonymous):

thank you hero, your not interrupting...and based on your formula i got 48419.032, but i do appreciate having someone explain the whole process to me which smoothmath is willing to do

OpenStudy (anonymous):

Wrong.

OpenStudy (anonymous):

wrong?

OpenStudy (anonymous):

Wrong.

OpenStudy (anonymous):

....which part the answer or the willing to explain

OpenStudy (anonymous):

If you want to do it Hero's way, let him explain it.

OpenStudy (anonymous):

woah now ill take all the help i can get!...if you want to keep explaining to me how to do it your way @SmoothMath then by all means continue, i always appreciate it when you teach me how to do this stuff it helps me a lot

OpenStudy (anonymous):

I think that good learning takes patience and focus. When other people are jumping in and trying to give you a shortcut or take you right to the answer without considering the process to get there, it hurts your understanding and slows things down.

OpenStudy (anonymous):

Maybe it's selfish or rude, but if I'm going to take my time and teach a student, I want to be their only teacher. Multiple teachers just doesn't work.

OpenStudy (anonymous):

ok well from this point on @SmoothMath shall be my only teacher for this problem...everyone else that comes in to contribute shall be shunned..and put to death!....better? i do understand your point and i think that patience and focus are attributes that every good teacher should possess...now are you still willing to walk me through how to do this? it would really help..i have a test tomorrow..

OpenStudy (anonymous):

Yes =) This is a very important skill in general. You are using a known formula, plugging in the variables you know, and then using algebra to solve for the unknown variable.

OpenStudy (anonymous):

Step 1: Figure out the right formula Step 2: Plug in the things you know Step 3: Solve for the thing you don't know

OpenStudy (anonymous):

reminds me of AP chem lol...my favorite...anyway continue

OpenStudy (anonymous):

Okay, so you said that you understand the step 1 here. We decided to use the formula \(\Large S_n = a_1\frac{1-r^n}{1-r}\) And hopefully we know what each of those variables means. If you don't, you should ask.

OpenStudy (anonymous):

well i dont know what the r represents

OpenStudy (anonymous):

Thankyou for saying so =) The idea with a geometric sequence is that to get to the next number, we multiply by some number. Here's an example of a really simple geometric sequence. 1, 2, 4, 8, 16, 32, 64 Do you see how it's doubling every time? So the number we multiply by to get the next number is 2. Now, that number multiply by is called the common ratio, and we usually give it the variable name r So for that sequence, r=2

OpenStudy (anonymous):

In your problem, r=3.2 So what that means is that I get the next number by multiplying by 3.2.

OpenStudy (anonymous):

o ok! well thats easy...continue

OpenStudy (anonymous):

Do you understand the difference between \(A_n \text{ and } S_n\)?

OpenStudy (anonymous):

an is nth term itself sn is sum of n terms...right?

OpenStudy (anonymous):

Right, exactly. So in my example sequence, 1,2,4,8,16 \(\large A_3\) would be 4 and \(\large S_3\) would be 7

OpenStudy (anonymous):

looks good!

OpenStudy (anonymous):

Okay, so we're good with step 1. Step 2 is to look at the info you know and plug into the formula. The formula we have is \(\Large S_n = A_1*\frac{1-r^n}{1-r}\) And the info we're given is \(S_n = 48455, r=3.2, n=3\)

OpenStudy (anonymous):

okay...so it would be S(3)=A(1)* 1-3.2^3/1-3.2

OpenStudy (anonymous):

Very good =) And we can do a little bit better than that, even. Since you're given Sn = 48455 when n=3.

OpenStudy (anonymous):

so just 48455=A(1)* 1-3.2^2/1-3.2?

OpenStudy (anonymous):

Exactly. Now the last step is just algebra, but this is the kind of algebra problem that a lot of students struggle with a little bit. Do you know how to solve for A1 here?

OpenStudy (anonymous):

well wishfully i would like to solve 1-3.2^2/1-3.2 and then divide my answer by both sides but im thinking thats not gonna fly..so im assuming i have to divide both sides by 1-3.2^2/1-3.2?...

OpenStudy (anonymous):

The first thing you said totally works. Do \(\Large \frac{1-3^2}{1-3}\). That'll make your problem a lot simpler.

OpenStudy (anonymous):

The second thing you said totally works too =) It's just not as simple, I'd say.

OpenStudy (anonymous):

Oh, oops. I said that wrong. Do \(\Large \frac{1-3.2^3}{1-3.2}\)

OpenStudy (anonymous):

lol well usually im all for the easy! but i just went ahead and did the long way...and i got 2383.66...is that right?

OpenStudy (anonymous):

You made a mistake somewhere, which is why I like the first way. You probably won't make the same mistake if you simplify it first.

OpenStudy (anonymous):

ok let me try again...one sec

OpenStudy (anonymous):

okay i did 1-3.2^3/1-3.2=-34.968...48455/-34.968= -1385.695493

OpenStudy (anonymous):

I think you're making a mistake in the order of operations somehow.

OpenStudy (anonymous):

Do the top and the bottom separately, then divide.

OpenStudy (anonymous):

Okay, yeah, I see what you did. You're probably putting it into your calculator without the right parenthesis.

OpenStudy (anonymous):

You got -34.968, which is what you get if you do \(\Large 1-\frac{3.2^2}{1}-3.2\)

OpenStudy (anonymous):

\(3.2^3\) I mean.

OpenStudy (anonymous):

Whenever you put a fraction into your calculator, you have to make sure you put the parenthesis around the correct things so that your calculator divides correctly.

OpenStudy (anonymous):

alrighty well i did what you said and i tried doing them separately and i got 14.44...then that divided by 48455 and got...

OpenStudy (anonymous):

14.44 is right =)

OpenStudy (anonymous):

355.609

OpenStudy (anonymous):

oops wait

OpenStudy (anonymous):

3355.609

OpenStudy (anonymous):

So you get 48455 = \(a_1*14.44\)

OpenStudy (anonymous):

Good =) Exactly.

OpenStudy (anonymous):

Okay, so I think that we can conclude that the part you need to work on the most in this whole thing is just the basic algebra.

OpenStudy (anonymous):

yes!! thank you so much!! this is WAY easier than i thought it was!!

OpenStudy (anonymous):

and ya i agree

OpenStudy (anonymous):

You're fine with setting up the formula, I think.

OpenStudy (anonymous):

You're welcome =)

OpenStudy (anonymous):

your going to be an amazing math teacher!!! thank you thank you thank you again! lol

hero (hero):

@SmoothMath, I never wanted to interrupt, but I was the one helping her originally. I had the formula wrong and I was trying to make up for it. I found the correct formula, and then you jumped in.

hero (hero):

You wouldn't allow me to make up for my mistake. That's what I don't appreciate.

OpenStudy (anonymous):

Yeah, well. I don't consider what you were doing to be an explanation. If you were teaching, I'd leave you be.

hero (hero):

@SmoothMath, I would have provided a better explanation, but you interfered.